OpenStudy (kainui):

@dan815 factoring a number "fourier series coefficient" style

OpenStudy (usukidoll):

I thought we can only factor term by term with integration for fourier series?!

OpenStudy (kainui):

I was thinking what if we had a number like this \[\Large \log n = \sum _{i=1}^\infty a_i \log p_i\] Is there some kind of thing we can do to solve for the individual coefficients like \[\Large \int\limits \log p_i \log p_j =0 \text{ for i} \ne \text{j}\]

OpenStudy (usukidoll):

this sounds similar to the proof I had to do last semester.

OpenStudy (dan815):

very interesting

OpenStudy (kainui):

Yeah in a way this is kind of like inverting a matrix and multiplying both sides of the equation by it so that we get the identity matrix

OpenStudy (dan815):

first does this integral really hold true?

OpenStudy (dan815):

for i=/=j

OpenStudy (kainui):

Dan no, it doesn't even have bounds. I'm saying let's find out what it has to be. We might have to include some other things, or possibly even make it a summation instead of an integral.

OpenStudy (usukidoll):

\[i \neq j \]

OpenStudy (kainui):

Like for instance: \[\Large f(x) = \sum_{n=1}^\infty a_n \sin(nx)\]\[\Large \int\limits_0^{2\pi} \sin(nx) \sin(mx) \frac{dx}{\pi} = \delta_{nm}\] Do you see the resemblance to matrix multiplication here by its inverse?

OpenStudy (dan815):

im thinking maybe, lets try to determine a relationship with logs such that we have an orthonormal basis to work with

OpenStudy (kainui):

it doesn't have to be orthogonal or normal, we can always orthogonalize it with gram-schmidt and normalize it as well later. It just has to form a basis that spans the set fortunately. =D

OpenStudy (dan815):

i dont understand gram-schmidt process for functions

OpenStudy (dan815):

does it come from the descrete method of solving for the orthgonal vectors

OpenStudy (kainui):

Well just treat the integral as the dot product and just do the same process you would normally for gram schmidt, just like projections. I actually just basically made a guide for this the other day haha

OpenStudy (dan815):

okay i think i can work that out there, its just like doing an integral in place of the dot products right and doing multiple integrals

OpenStudy (dan815):

ill take a look at it after this

OpenStudy (dan815):